# Mass measurements in the 132Sn region with the JYFLTRAP double Penning trap mass spectrometer

### v2

@@ III Results and discussion
An updated mass value of 137Sb leads to a more precise neutron-capture $Q$-value, $Q_{n,\gamma}=3576(9)$ keV. Its impact on the neutron-capture 136Sb$(n,\gamma)^{137}$b reaction rate was studied using the Talys 1.96 code [55]. A comparison between this work (JYFLTRAP), AME20 [22], and ESR [17] is presented in Fig. 5. The reaction rate obtained with the 137Sb mass value from this work and AME20 agree with each other. However, at a typical r-process temperature of $T=1$ GK the former is around 7% lower than the latter and, in addition, it is around five times more precise. The reaction rate calculated with the ESR result significantly deviates from both, this work and AME20, and it is about twice larger.

### v3

@@ III Results and discussion
Since for the iodine isotopes the determined mass values are either in agreement with, or very close to, the literature values (see Table 1), the effect of the new masses on the calculated neutron-capture rates for the r process is almost negligible as compared to AME20 and is therefore not discussed further. The biggest difference is observed for 137Sb where the updated mass value leads also to a more precise neutron-capture $Q$-value, $Q_{n,\gamma}=3576(9)$ keV, as compared to the AME20 value, $Q_{n,\gamma}=3620(50)$ keV [22]. Our $Q$-value rules out the value based on the 137Sb mass from the ESR measurement, $Q_{n,\gamma}=4170(170)$ keV.
The impact on the neutron-capture 136Sb$(n,\gamma)^{137}$b reaction rate was studied using the Talys 1.96 code [56]. A comparison between this work (JYFLTRAP), AME20 [22], and ESR [17] is presented in Fig. 6 for temperatures $T\leq 5$ GK relevant for neutron captures in the r process (see e.g. Refs. [2, 57]). At higher temperatures, photodisintegrations dominate and abundances are determined by the nuclear statistical equilibrium. The reaction rate obtained with the 137Sb mass value from this work and AME20 agree with each other. However, for example at a characteristic r-process temperature of $T=1$ GK [57], the former is around 7% lower than the latter and, in addition, it is around five times more precise. The reaction rate calculated with the ESR result significantly deviates from both, this work and AME20, and it is about twice larger. Thus, our measurement rules out such a neutron-capture reaction rate.